### v1

Abstract
Context:The kinematics and chemical composition of stellar populations of different ages provide crucial information on the evolution of the various components of a galaxy.
Aims:Our aim is to determine the kinematics of individual stars as a function of age in IC 1613, a star-forming, gas-rich and isolated dwarf galaxy of the Local Group.
Methods:We present results of a new spectroscopic survey of IC 1613 conducted with MUSE, an integral field spectrograph mounted on the Very Large Telescope. We extracted sources, from which we separated stellar objects for their subsequent spectral analysis. The quality of the dataset allowed us to obtain accurate classifications ( to better than K) and line-of-sight velocities (with average km s) for about 800 stars. Our sample includes not only red giant branch (RGB) and main sequence (MS) stars, but also a number of probable Be and C stars. We also obtained reliable metallicities ( dex) for about 300 RGB stars.
Results:The kinematic analysis revealed for the first time the presence of stellar rotation with high significance. We found general agreement with the velocity field of the neutral gas component, although on average the stars showed greater velocity dispersion and slower rotation, which can be attributed to asymmetric drift. Examining the kinematics of the different stellar components, the MS stars appear to follow closely that of the gas. The chemical analysis of the RGB stars revealed average properties comparable to other Local Group dwarf galaxies.
Conclusions:Our work represents a step forward in understanding the internal processes that govern the dynamical evolution of a low-mass galaxy.
From gas to stars: MUSEings on the internal evolution of IC 1613
1 Introduction
The ability to resolve stellar populations in nearby galaxies allows us to study the physical processes that govern their evolution. The dwarf galaxies of the Local Group (LG), in particular, provide an excellent laboratory due to their diverse characteristics, including size, gas content, star formation history (SFH), chemical abundances, and kinematics (e.g. Tolstoy et al. 2009; McConnachie 2012; Kirby et al. 2013; Gallart et al. 2015; Simon 2019; Putman et al. 2021; Battaglia & Nipoti 2022). Since these characteristics can be influenced by the proximity of a larger galaxy, such as the Milky Way and Andromeda, isolated LG dwarf galaxies provide valuable insights into the intrinsic mechanisms that regulate their evolution. In the gas-rich ones, the interplay between instability and stellar feedback processes can be reconstructed by combining the kinematics and spatial distribution of stars of different ages with the SFH and gas content of the galaxy (e.g. Leaman et al. 2017; Collins & Read 2022). Gathering this kind of information is usually an observational challenge, but the advent of integral field spectroscopy (IFS) has been transformative in this respect, due to the ability to inspect different galactic tracers with high observational efficiency (Roth et al. 2018, Sánchez et al. 2012; Cortese et al. 2014; Bundy et al. 2015; McLeod et al. 2020; Zoutendijk et al. 2020; Júlio et al. 2023; Vaz et al. 2023, but also).
Among LG dwarf galaxies, IC 1613 is an ideal IFS target. On the one hand, it is a typical low-mass star-forming dwarf irregular galaxy (Skillman et al. 2014). On the other hand, stellar feedback processes have generated voids and bubbles in the interstellar medium whose impact on stellar and gas kinematics needs to be explored (Read et al. 2016). First discovered by Wolf (1906), IC 1613 was recognised soon afterwards as an extragalactic object by Baade (1935, as reviewed by Sandage 1971), who first determined its distance using Cepheid variable stars. Since then, several other measurements of its distance have been made using different indicators, including RR-Lyrae variable stars and the tip of the red giant branch (RGB). A compilation of literature distance values (including their own) was reported by Bernard et al. (2010), who obtained a statistical average value of , or kpc.
IC 1613 therefore lies well beyond the Milky Way’s (MW) virial radius, as well as that of M31 (being at kpc distance from it, McConnachie 2012). The systemic proper motion of IC 1613 obtained from McConnachie et al. (2021) with Gaia DR2 data does not rule out the possibility that the galaxy was ever within 300 kpc of M31, for M31 masses of . However, the most recent estimate of the systemic proper motion of IC 1613 by Bennet et al. (2023), combining data from Gaia eDR3 and the Hubble Space Telescope (HST), makes such an association unlikely when assuming a mass of M31 of . Interestingly, Buck et al. (2019) assign to IC 1613 a high probability of being a backsplash galaxy (i.e. a currently isolated system that may have once passed close to a large host, but without becoming bound) of M31.
Its low luminosity (an absolute magnitude of ) corresponds to a stellar mass of , similar to that of other LG dIrrs like NGC 6822 and WLM (McConnachie 2012). The SFH and stellar content of IC 1613 have been studied in detail over the years (see, e.g., Cole et al. 1999; Skillman et al. 2003; Bernard et al. 2007; Skillman et al. 2014). Based on deep HST Advanced Camera for Surveys (ACS) imaging, and after a review of archival HST data, Skillman et al. (2014) concluded that the SFH of IC 1613 has been close to constant on average throughout its life, with no evidence of an early dominant episode of star formation. We note that the HST/ACS field analysed by Skillman et al. (2014), roughly located at the half-light radius of the galaxy, was considered as representative of the global SFH of IC 1613. The comparison with archival HST observations located at different radii supported this assumption. The average metallicity was linearly increasing through time, with values ranging from dex to dex. Stars were formed at an average rate of .
The structural properties of IC 1613 have been determined by studying the spatial distribution of stellar tracers in different evolutionary phases (see, e.g., Albert et al. 2000; Borissova et al. 2000; Bernard et al. 2007; Garcia et al. 2009; Sibbons et al. 2015; McQuinn et al. 2017; Pucha et al. 2019; Higgs et al. 2021). In general, it was found that the young stars are more centrally concentrated than the intermediate-age and old stars, a feature that is common in dwarfs. However, Pucha et al. (2019) using deep and wide Subaru/Hyper-SuprimeCam observations of IC 1613, showed that its young main sequence (MS) stars, along with the RGB and ancient horizontal branch (HB) stars, all extend to the outskirts of the galaxy, up to arcmin (i.e., effective radii ). In particular, the young stars are found well beyond the currently active star forming regions (within ), although with a much lower density compared to the intermediate and old age components. They also showed a steeper radial density profile in the inner regions than in the outer ones, in contrast to the RGB and HB stars. This seems to imply a different formation channel between the younger (e.g., from gas pushed outward through stellar feedback) and older stars (probably via gas accretion) in the galaxy’s outskirts. In this work, we assume the structural parameters from Higgs et al. (2021) who conducted an homogeneous analysis for the isolated LG dwarf galaxies.
As other dIrrs, IC 1613 contains an extended component of neutral hydrogen (HI) gas, with a clumpy distribution on the inside, but showing regular contours at larger radii (see van den Bergh 2000, and references therein). In particular, the HI distribution is rich of shells and voids around the currently active star forming regions, where also the ionized gas is distributed (e.g., Lozinskaya et al. 2003; Silich et al. 2006; Moiseev & Lozinskaya 2012; Pokhrel et al. 2020). The HI kinematics shows a linearly increasing rotation curve with a km s (Lake & Skillman 1989; Oh et al. 2015; Read et al. 2016). The rotation value, however, may represent a lower limit due to the fact that this galaxy is probably seen face-on. This translates into a significant uncertainty in the determination of the dynamical mass being between and , depending on the assumed inclination angle (Read et al. 2016).
The first spectroscopic measurements of stars in IC 1613 were obtained from some of the brightest ones, that is, from 9 early B-type young supergiants (Bresolin et al. 2007) and 3 evolved M-type supergiants (Tautvaišienė et al. 2007). Both studies obtained compatible metallicity values, dex and dex, respectively111Assuming dex and a solar oxygen abundance of dex (Asplund et al. 2009)., later confirmed by Berger et al. (2018) who studied 21 young BA-type supergiant stars. Interestingly, these authors found a bimodal metallicity distribution that appears to be correlated with their spatial location (i.e. they have lower metallicity in regions of high HI density). A comparable bi-modal distribution was also observed by Chun et al. (2022) for a sample of 14 red supergiants. Furthermore, there is evidence from the youngest stellar population and evolved red supergiants of IC 1613 that the present-day ratio could be sub-solar (i.e., dex, Tautvaišienė et al. 2007 and Garcia et al. 2014). The chemical abundance of the interstellar medium has been obtained spectroscopically from the central HII regions and is in good agreement with results from the supergiant stars (Bresolin et al. 2007, but see also Lee et al. 2003 and Tautvaišienė et al. 2007).
Kirby et al. (2013) were the first to present a statistically significant spectroscopic analysis of a sample of 125 RGB stars in IC 1613, which led to an average metallicity of dex, in agreement with the fact that older stars are in general more metal-poor than the younger ones, and with expectations from the age-metallicity relation found by Skillman et al. (2014). On the same sample, Kirby et al. (2014) conducted a kinematic analysis which led to a systemic velocity of km s, in agreement with values from the HI component (Lake & Skillman 1989; Oh et al. 2015), and a velocity dispersion of km s. They did not find signs of rotation for the stellar component; this was later confirmed also by Wheeler et al. (2017) who re-analysed the Kirby et al. dataset. This is not surprising considering that the spatial distribution of this dataset is roughly along the optical minor axis of the galaxy, while in general it is expected that the stellar rotation follows that of the gas, which in this case rotates roughly along the optical major axis (Oh et al. 2015). The dynamical mass at the half-light radius reported by Kirby et al. (2014) was of .
In this paper, we present a study of the stellar kinematic and chemical properties of IC 1613 from spectroscopic data taken with the MUSE integral field spectrograph on the Very Large Telescope (VLT). Taking advantage of the unique capabilities of this instrument, that combines high spatial resolution with a 1 arcmin field-of-view and a wide wavelength range at medium spectral resolution (Bacon et al. 2014), we performed a kinematic analysis as a function of stellar age and obtained metallicites for the largest spectroscopic stellar sample obtained to date for this galaxy. IC 1613 is located at high Galactic latitude ( deg), so its low values for both foreground (Schlafly & Finkbeiner 2011) and internal reddening (Georgiev et al. 1999) make it an ideal laboratory to compare the observed properties of its diverse stellar content.
The article is structured as follows: in Sect. 2 we give details on the data acquisition and reduction processes we conducted. Section 3 is dedicated to spectral classification and velocity determination. In Sect. 4 we report details on the determination of likely member stars and their subsequent kinematic analysis. Section 5 shows results of the chemical analysis of the RGB stars. Finally, Sect. 6 is dedicated to the summary and perspectives of future work, while in the appendices we report details on the sanity checks conducted on the spectral-fitting codes used during our analysis. The parameters adopted for IC 1613 throughout the text are summarized in Table 1.
| Parameter | Units | Value | Ref. |
| kpc | (2) | ||
| arcmin | (3) | ||
| kpc | (3) | ||
| P. A. | deg | (3) | |
| 0.200.05 | (3) | ||
| P. A. | deg | (4) | |
| deg | (4) | ||
| km s | (5) | ||
| km s | (5) | ||
| dex | (5) | ||
| dex | 0.26 | (5) | |
| Conversion | pc arcmin | 221 |
References. (1) McConnachie (2012); (2) Bernard et al. (2010); (3) Higgs et al. (2021); (4) Read et al. (2016); (5) this work.
2 Observations and Data Reduction
The data were acquired with VLT/MUSE in service mode333Under ESO programme 097.B-0373; PI: Battaglia.. MUSE is an integral field spectrograph with a spatial sampling of 0.2 arcsec and a spectral resolution varying between along its wavelength coverage. We used the instrument in nominal wide field mode, with a field of view of and a wavelength coverage between Å. We observed 3 fields (hereafter F1, F2, F3, moving outward from the centre; see Fig. 1 for their location) approximately along the projected semi-major axis of IC 1613, on the West side of the galaxy. We purposely avoided regions of on-going star formation.
Each field was observed with a total exposure time of 11080 s on source, split in s exposures, obtained by rotating the position angle of the spectrograph by 90 with respect to the previous exposure. No separate sky field was acquired since the stellar component of IC 1613 is resolved in the magnitude range of interest and the sky contribution is taken care of during the de-blending and spectral extraction phase of the data analysis (see Sect. 2.1). The data were acquired in dark time, clear sky, with a request for a seeing in V-band at zenith arcsec, fulfilled for all but one exposure of F1 (for which it deviated within 10% from the request). The observing log is reported in Table 7.
2.1 Data Reduction and Source Extraction
All data cubes were reduced using version 1.6 of the official MUSE pipeline (Weilbacher et al. 2012). The basic steps of the reduction cascade (bias subtraction, slice tracing, wavelength calibration, and basic processing) were performed using the default settings444This included correcting for the observer’s line of sight velocity with respect to the barycentre of the solar system.. This resulted in 24 pixel tables (one for each unit spectrograph) for each individual exposure. When combining the pixel tables for each exposure, flux calibration was also applied using standard stars observed on the same nights. Furthermore, the subtraction of sky emission lines was performed by determining their intensity directly from the scientific data, which contained sufficiently large patches of (almost) blank sky for this purpose. The sky emission lines were also used to quantify the quality of the wavelength calibration, which had an average accuracy of 0.03 Å, or 1.5 km s. The final step of the data reduction was the combination of all individual exposures obtained for a pointing. The end products of the data reduction were three data cubes, one for each field. Each cube has a dimension of approximately pixels.
The PampelMUSE software described in Kamann et al. (2013) was used to extract the stellar spectra from the reduced data cubes. To run PampelMUSE, an input catalogue of source positions and an estimate of their initial magnitudes is required. Since the MUSE data are only moderately crowded, the input catalogues were directly created from the data cubes. We produced synthetic broad-band V-, R- and I-band as well as emission line color-composite images (see Fig. 2), using the procedures of Roth et al. (2018), and extracted photometric catalogues using the DAOPHOT code (Stetson 1987). The photometric catalogues were obtained as in Gallart et al. (1996). They proved to be a valuable resource for our analysis, as they contain the spatial and photometric information of all the extracted sources. The details of their astrometric and photometric calibration are reported in Appendix A.
The raw source catalogues (i.e. without astrometric solution or photometric calibration applied) were passed as input to PampelMUSE to first identify the positions of stars in the data cubes as a function of wavelength. The code then aimed to retrieve the point spread function (PSF) using an analytical Moffat profile. To obtain the final spectra, a simultaneous PSF fit was performed to all sources in each layer of the data cube; unresolved components were treated by including a background in the fit that was recalculated at fixed spatial offsets. As IC 1613 contains gaseous emission that varies on rather small spatial scales, well visible in Fig. 2, the background component was recomputed every 40 spaxels.
PampelMUSE extracted 2293 initial spectra, examples of which are shown in Figs. 11, 12 and 13. For each extracted spectrum, PampelMUSE provided the associated flux uncertainty per pixel and an S/N value calculated around its central wavelength (hereafter S/N). However, as our sources cover a wide spectral range (see Sect. 3.1), we calculated two further S/N indicators: one around 5500 Å (S/N), and the other on the continuum around the CaT lines (S/N). This gave us a more accurate picture of the quality of our spectra depending on the spectral type of a given star.
The photometric catalogues were used to perform a first cleaning of the spectroscopic sample, excluding clearly non-stellar sources. We selected targets based on specific photometric parameters, including sharpness between and , and a goodness-of-fit (CHI) of less than 1. We also excluded targets with magnitudes in only one band, which were generally low S/N targets identified only in the I band. As a result, the total number of extracted sources decreased from 2293 to 2053.
The quality of our data can be seen in Fig. 3. In particular, we obtained respectively in F1, F2, F3 about 65, 40, 10 sources with S/N, and about 200, 120, 45 sources with S/N. We note that for this galaxy dust reddening and extinction along the line of sight are almost negligible, with average values of and (Schlafly & Finkbeiner 2011). Nevertheless, we have taken this into account.
3 Spectral types and velocity determination
We performed an analysis to determine the spectral type and line-of-sight (l.o.s.) velocity of our sources. For this purpose, we used, respectively, the spectral fitting codes ULySS (Koleva et al. 2009) and spexxy (Husser 2012).
3.1 Spectral classification with ULySS
The spectral type classification was carried out with the same approach as in Roth et al. (2018); we refer the reader to this article for full details. A linear combination of up to 10 templates from the MIUSCAT stellar spectral library555The MIUSCAT stellar spectral library fills in the gap between the MILES (Sánchez-Blázquez et al. 2006) and the near-IR CaT library of Cenarro et al. (2001), using the Indo-US spectral library of Valdes et al. (2004); it also extends the wavelength coverage slightly blue-ward of MILES and red-ward of the CaT library. MIUSCAT has been recently replaced by E-MILES (Vazdekis et al. 2016). (Vazdekis et al. 2012) was fitted to the observed spectra using the ULySS666http://ulyss.univ-lyon1.fr code (Koleva et al. 2009). We note that ULySS is capable of identifying unresolved blended stars because the spectrum is fitted with a linear combination of non-linear components, multiplied by a polynomial continuum. In our case, no blends are found, not even in the central field. This is mostly expected, given the density of sources visible in Fig. 2, which was constructed from the MUSE data cubes.
We stored the information of the three spectral types having the largest weights in the best fit. Occasionally, the fit returns implausible mixes of spectral types (e.g., mixing young early-types with evolved late-types). The apparent magnitude obtained by shifting the absolute magnitude in R-band of the stars in the MIUSCAT library to the distance of IC 1613 (distance modulus , from Bernard et al. 2010), provided a way to weed out these clearly erroneous fits when they would lead to apparent magnitudes well below the detection limit of the MUSE data. However, we also found that good-quality fits to cool MS stars allowed us to confidently identify Milky Way foreground stars that would otherwise have been difficult to reject from photometry alone.
The ULySS spectral analysis was carried out on all those sources having a S/N and a clean photometry (see Sect. 2.1). Main outputs of the code were the best fitting templates, together with the associated spectral parameters (, log(g) and [Fe/H]) and l.o.s. velocity. We verified by eye the ULySS outcomes, assigning several quality flags to each analysed object. We evaluated the quality of the input spectrum (QSP), the quality of the spectral fitting (QFT), the plausibility of the output l.o.s. velocity (PVR) and the plausibility of the spectral type classification (PCL); flags were reported as integer numbers ranging from one to four, with the lower value meaning a useless measure, and the higher value implying an excellent fit. In this work, we used ULySS exclusively for spectral classification purposes, while we left the l.o.s. velocity determination part to the spexxy code. This was mainly because spexxy performed better than ULySS in the velocity determination task, especially in the low S/N regime (i.e., ), as we verified in Appendix B.1 and B.2.
| Sp. Type | (K) | N |
| M | 1 | |
| K | 467 | |
| G | 90 | |
| F | 46 | |
| A | 57 | |
| B | 96 | |
| O | 13 | |
| Be | 24 | |
| C | 14 | |
| Total | 808 |
In Table 2, we report results from the spectral classification obtained for sources marked with a PCL flag between three and four (which implies an accuracy on the effective temperatures better than 500 K). The visual inspection during the spectral classification also allowed us to identify other contaminants that were not removed during the previous photometric selection (Sect. 2.1): they were mainly background galaxies (comprising high-z emitters) and low S/N targets highly polluted by diffuse ionised gas emission lines. We also visually identified carbon and B emission line stars that were missed by ULySS due to the lack of adequate stellar templates, in order to study them separately (see Sect. 3.3). We note that we define as bona-fide B emission-line stars (noted as Be hereafter) those sources whose spectral distribution is compatible with a star having K, but which show H and (occasionally) H in strong emission (e.g., Porter & Rivinius 2003; Rivinius et al. 2013). The other seven stars showed only H in weak emission, but due to their low S/N their classifications were too uncertain and they were therefore discarded.
Overall, we obtained a reliable spectral classification for 808 stars888The spectroscopic for the red giant stars were found to be in fair agreement with the empirical relation from Alonso et al. (1999), which is valid in the colour range and mostly metallicity independent.. Looking at Table 2, the large majority of sources are K-type giants, with an almost even representation of the other types, except for the M and O stars being a minority. In particular, among the M stars, only one was found to be a giant, while the other four were classified as MW-foreground dwarfs. We also report the presence of two other MW-contaminants classified as K V and G V stars, all confirmed in the l.o.s. velocity determination step. We have therefore excluded these MW-foreground stars from the counts in Table 2 and the following analysis. Fig. 4 shows the position of the classified stars on the colour-magnitude diagram, colour-coded according to their , visually confirming the general goodness of the spectral classification.
The number of O-type stars identified is the most uncertain, since the poor coverage of the MIUSCAT library at the highest effective temperatures (i.e. for K) is a limiting factor for the spectral classification of these stars. We have identified 13 stars with a just above 30000 K. This is in general agreement with the expected number of O stars in IC 1613. Indeed, considering the galaxy’s current star formation rate (Skillman et al. 2014) and assuming a Kroupa (2001) initial mass function, we would expect on the order of ten O stars in the surveyed area, given their average lifetime of 10 Myr and minimum mass of .
On the other hand, O stars are typically not uniformly distributed in space. They tend to form and evolve in OB associations. Our observations cover some of the OB associations identified by Garcia et al. (2009). These associations are low density groups of young stars with significant internal extinction, making the identification of O types more challenging. In Fig. 5 we show the spatial distribution of the OB associations reported by Garcia et al. (2009), together with that of our identified OB stars with K and also the Be stars. We see that they generally tend to be found where the OB associations are. Their spatial distribution also coincides with the ionised shells shown in Fig. 2, to which they are probably physically associated (Borissova et al. 2004; Garcia et al. 2009). At this stage, these evidences at least confirm the goodness of our initial classification of many of them as hot stars.
We further note that the identified Be stars account for up to 18% of the total OB sample. This is comparable to the observed Be fraction in nearby metal-poor dwarf galaxies (between 15% and 30%, Schootemeijer et al. 2022; Gull et al. 2022; Vaz et al. 2023). Since Be stars are likely to be highly rotating massive stars, our results add to the evidence that they are common in low-metallicity environments.
We are currently preparing a follow-up study using a grid of model atmospheres obtained from the NLTE code FASTWIND (Puls et al. 2005), whose spectra will allow us to fit the more massive stars of spectral type AO that are sparsely covered by the empirical MIUSCAT library. Nevertheless, the wavelength range of MUSE, which does not cover the gravity and temperature sensitive lines below 4800 Å, is a limitation for the analysis of massive stars. An accurate determination of these parameters, as well as Fe and -elements abundances, will be possible with the future BlueMUSE instrument (Richard et al. 2019).
Looking at the coolest stars in our sample, we have identified a single M-type giant and 14 carbon (C) stars. The latter are a type of asymptotic giant branch (AGB) stars, characterised by having more carbon than oxygen in their atmospheres. This happens during the thermally pulsating (TP) phase, when C-rich material from the star’s interior is dredged up into the previously oxygen-rich atmosphere (a process known as the third dredge-up). A TP-AGB star is initially a O-rich M-type giant. Since it is easier to transition from an O-rich to a C-rich star when its metallicity is low, the C/M fraction is usually an indirect [Fe/H] indicator of intermediate-age stars in a galaxy. For IC 1613, literature works report an average value of C/M (see e.g. Albert et al. 2000; Chun et al. 2015; Sibbons et al. 2015; Ren et al. 2022, based on wide-area JHK-photometric surveys).
We can qualitatively determine the C/M in our spectroscopic sample. The only M-type giant was clearly an O-rich AGB star. We also examined the K-type stars above the tip of the RGB and selected a sample with within 500 K of an early M-type star. Cross-correlating with the literature photometric catalogues (Albert et al. 2000; Chun et al. 2015; Sibbons et al. 2015), we found 10 common targets classified as M stars (after excluding the possible red supergiants identified by Ren et al. 2022). As for our C stars, three have already been identified in the literature (see again Albert et al. 2000; Chun et al. 2015; Sibbons et al. 2015), while five have a more uncertain spectral classification due to their faint magnitudes (), well below the tip of the RGB. It should be noted that TP-AGBs are long-period variables (with periods of a few tens to hundreds of days), which can exhibit brightness excursions of several magnitudes (see e.g., Menzies et al. 2015). Alternatively, faint C stars can be binary systems in which the primary star has received C-enriched material from the secondary star, which was previously an AGB (also known as extrinsic C stars). For IC 1613, the expected fraction of extrinsic C stars is rather low (, Hamren et al. 2016, e.g.,). Considering between nine and 14 C stars, the C/M ratio then ranges between 0.9 and 1.4, which is larger than the average value reported in the literature. However, taking into account the Poisson error associated with the small number statistics of our sample, we would remain in agreement with the literature.
3.2 Velocity determination with spexxy
The determination of l.o.s. velocities was carried out using the spexxy999https://github.com/thusser/spexxy code (v. 2.5, Husser 2012). The routine performs a full spectral fitting to the observed spectra using interpolated spectral templates generated from the PHOENIX library of high-resolution synthetic spectra101010http://phoenix.astro.physik.uni-goettingen.de/ (Husser et al. 2013). The library covers a wide wavelength range, going from 500 Å to 5.5 m, and stellar parameters: , , and . However, the PHOENIX models do not treat radiative transfer in the atmospheres of massive stars affected by strong stellar winds, so the library is of little use in the recovery of stellar parameters for K. We refer to Appendix B.1 for a detailed explanation of the spectral fitting steps and to B.2 for a comparison of its performance with the ULySS results presented above.
We ran spexxy on all targets having S/N (in at least one of the S/N indicators) and clean photometry (see Sect. 2.1). We also excluded all those targets that in the spectral classification step (see Sect. 3.1) were identified as contaminated by ionised gas, background galaxies, and Be stars (which we analysed separately). These steps reduced our sample to 1745 input sources. We were thus able to recover their l.o.s. velocity, along with estimates of their spectral parameters and associated errors. We note that only for of stars spexxy failed to perform a successful fitting, mostly due to their low S/N.
The goodness of the recovered parameters was quantified in a series of tests, whose details are given in Appendix B.1. Here we report the essentials. We verified by performing several tests on mock spectra that the l.o.s. velocity errors are well estimated down to S/N for red giant stars (S/N for MS stars), and the velocity values are recovered without significant offsets. On the other hand, the recovery of the spectral parameters was more limited, with in particular being generally well constrained for red giants, but underestimated for the hottest MS stars due to the grid limitation of our spectral models. We emphasise again that our aim was to use spexxy mainly for the determination of l.o.s. velocities.
Since our main goal is to analyse not only the global kinematic properties of IC 1613, but also how they change as a function of stellar types, we decided to keep the sample obtained during the spectral classification step. From this sample we further removed those targets with a velocity error km s, which meant excluding of sources with the least reliable velocity values (mostly low S/N MS stars). We also excluded the six targets previously classified as MW contaminants, confirming that they are foreground sources based on their l.o.s. velocities, which were consistent with 0 km s. Performing a cross-correlation with the third data release of the Gaia catalogue (Gaia Collaboration et al. 2023), we further confirmed their foreground nature based on their proper motion values, significantly different from 0 km s in this case. With these conditions applied, our sample reduced to 727 sources to which we subsequently added the 24 Be stars whose velocity determination we describe in the next section. The distribution of l.o.s. velocity measurements for our sample is shown in Fig. 6, with histograms divided by pointing and stellar type. We note that the median velocity error was of km s, while the median S/N resulted around 10.
3.3 Velocity determination for carbon and Be stars
The l.o.s. velocities of the C stars identified by spectral classification (see Sect. 3.1) were calculated using spexxy as for the main sample. However, the code was not always able to match the complex spectral features of these stars. Therefore, we double-checked the spexxy velocity determination with ULySS using stellar templates from the X-Shooter library of carbon stars (Gonneau et al. 2016). We found a general agreement within the errors, confirming the goodness of the spexxy velocities. The median velocity error of the 14 C stars analysed (6 in both F1 and F2, and 2 in F3) was km s for a median S/N.
On the other hand, the velocity determination for the Be stars was done separately, since spexxy failed in this task because such stars are not represented by the PHOENIX model atmospheres. We thus performed a cross-correlation with custom-made template spectra using the function correlation of the python package specutils, part of the Astropy project (Astropy Collaboration et al. 2022). Templates were generated as ad-hoc proxies using the spectra of two B-type stars (classified as B9III and B3III) from the MIUSCAT library (Vazdekis et al. 2012), to which we added two Gaussian profiles reproducing the strength and width of the H and H emission lines as detected in our sample. Since the spectra of our Be-star candidates in general did not show any split line profiles (except for two objects), the two-component Gaussian approximation seemed to be a good enough approach for the purpose of l.o.s. velocity measurements. We cross-correlated the observed spectra with the templates and assigned them the average of the derived velocities. The errors were instead assigned by a Monte Carlo process. For each template we generated mock spectra at different S/N (calculated around 5500 Å) values . For each S/N, we cross-correlated 250 mock spectra with the corresponding noise-free template and considered the median absolute deviation of the velocity distribution as the velocity error representative of that bin. The error distribution as a function of S/N was fitted by an exponential profile for each template and we used the mean profile to assign velocity errors to the observed spectra according to their S/N. For the 24 Be stars analysed (14, 6 and 4 in fields F1, F2 and F3, respectively), we recovered a median km s for a median S/N.
We note that if we perform the kinematic analysis described in the next section on both the Be and MS star samples, we find no significant differences between them in terms of systemic velocity and velocity dispersion. Therefore, we included them in the main sample when we performed the kinematic analysis.
4 Internal kinematics
| Sample | N | N | Model | k | Bayes factor | ||
| (km s) | (km s) | (km s arcmin) | ln() | ||||
| Main | 751 | 746 | Linear rotation (HI) | 4.5 | |||
| Linear rotation (optical) | 3.5 | ||||||
| Dispersion only |
We investigate the basic internal kinematic properties of the stellar component of IC 1613, such as its systemic l.o.s. velocity, its velocity dispersion and the possible presence of velocity gradients, indicating rotation of the stellar component121212On the angular scales covered by the MUSE sample, we find the perspective rotation due to the relative motion between the galaxy and the Sun to play a negligible role with up to km s (following Eq. 6 of van de Ven et al. 2006 and assuming the systemic proper motion from Battaglia et al. 2022). Velocity gradients induced by tidal disturbances from the MW or M31 are also highly unlikely..
To proceed with the kinematic analysis, we first need to identify the possible contaminants left in our sample. We had already removed many of them during the spectral classification process, with the rest expected to be faint foreground stars. We used then the method outlined in Taibi et al. (2020), which allows a Bayesian kinematic analysis while assigning membership probabilities to the individual targets. This approach is based on the expectation maximization technique presented in Walker et al. (2009).
The values depend on the l.o.s. velocities of the individual targets, but also on the prior information based on their radial distances from the galaxy’s centre and on the expected l.o.s. velocity distribution of the possible contaminants. The spatial prior takes into account that the probability of membership is higher, the closer to the galaxy’s centre. In this case, we simply required that our targets should follow a monotonically decreasing radial density profile applying an isotonic regression model, with no assumptions about its functional form131313https://scikit-learn.org/stable/modules/isotonic.html.
We used the Besançon model of Galactic foreground stars (Robin et al. 2003) to obtain the expected l.o.s. velocity distribution of the contaminants. We generated a catalogue along IC 1613’s direction, over an area up to its half-light radius, covering the range of colours and magnitudes of our targets. This resulted in a velocity distribution that was well sampled and representative of the spatial area around the MUSE pointings. We approximated it as the sum of two Gaussian profiles (with means km s, km s, and standard deviations km s, km s, with an amplitude ratio of ). According to this model, considering the small area covered by the MUSE pointings and the spanned range of magnitudes and colours, we expect to find, in total, around five foreground contaminants in the velocity range between km s and 200 km s. If we restrict to the velocity range covered by the input targets (i.e. between km s and km s, see Fig. 6), we would expect only one contaminant. We recall that already during the spectral classification step (see Sect. 3.1) we found and removed six likely contaminants, all being cold MS stars with velocities around .
We carry out a Bayesian analysis to explore and compare different kinematic models, considering one where the internal kinematics of IC 1613’s stellar component can be purely described by random motions, and another model which also contains a rotational term. For the rotational term we used the form: , where k is the velocity gradient, are the angular distance from the galaxy’s centre (i.e. , where () are the tangent plane’s coordinates) and position angle (P. A.) of the i-target, and is the P. A. of the kinematic major axis (both P. A. measured from North to East).
We note that, unless otherwise stated, for the stellar component we quote the observed rotational velocity and not its intrinsic value , as the latter requires knowledge of the inclination angle i of the angular momentum vector with respect to the observer (i.e., ). The stellar component of Local Group dwarf galaxies is not in a thin disc, therefore its inclination cannot be determined without knowledge of the intrinsic 3D shape. In those systems that contain a clearly rotating HI component, as IC 1613, one can at least assume the inclination determined from the neutral gas component, under the assumptions that stars and gas share the same inclination. This is what we will do when quoting values for or when deriving estimates for the circular velocity. However, we note the possibility that IC 1613 may be seen close to face-on (Read et al. 2016), therefore even the HI value has its limitations.
The particular spatial distribution of the MUSE pointings and the limited area they cover prevent us from finding . Therefore, we fixed the central coordinates and of the kinematic field to be either that of the stellar component or that of the HI gas (see values in Table 1). The free parameters were then the systemic l.o.s. velocity , the l.o.s. velocity dispersion common to both models, and the linear gradient k. We assumed the following prior ranges: , where is the mean value of the input velocity distribution, , . The analysis also returned the model evidence Z which resulted useful for comparing the statistical significance of one model against another through the use of the Bayes factor . We obtained the kinematic parameters and the membership probabilities using the MultiNest code (Feroz et al. 2009; Buchner et al. 2014), a multi-modal nested sampling algorithm.
Results of the probability-weighted analysis are reported in Table 3. Of the initial 751 targets, 746 have a . We found that rotation is strongly favoured over a dispersion-only model, with a higher evidence for the case that assumes the central coordinates and orientation of the kinematic major axis from the HI. In this case, the systemic l.o.s. heliocentric velocity of the model including a rotation term is km s, the velocity dispersion km s, and the velocity gradient was km s arcmin (see Fig. 7). As these values have the highest evidence, we adopt these as our reference values in the following analysis.
The systemic velocity and velocity dispersion values we found are within 1- from those previously published by Kirby et al. (2014, i.e. km s and km s, respectively). On the other hand, Wheeler et al. (2017), analyzing the data by Kirby et al. (2014), reported a rotation amplitude for IC 1613 which was largely unconstrained and consistent with null. This is not surprising as their data are spatially distributed roughly perpendicularly to the direction of our MUSE data-set, for which we do detect a clear rotation signal. Therefore, this is the first time that rotation of the stellar component is detected with high significance in IC 1613.
We note that the five low probability targets could still belong to IC 1613 based on their spectral type. Only one has , but we checked that the velocity determination in this case was affected by the presence of residual emission lines from the surrounding ionised gas. By avoiding these lines, we recalculated its velocity, which was compatible with , although with too large an error ( km s). The other stars have . Three of these are low S/N young stars with large velocity errors. Interestingly, the remaining star has and km s, which puts it more than 3 away from the expectation of the best fitting rotation model. However, it is found in F1 and has photometric properties and metallicity (see Sect. 5) in common with the RGB stars of the galaxy. Given the low probability of being a MW contaminant, this star could be a possible binary, but we cannot exclude that it could belong in projection to an extended hot stellar halo around IC 1613. Indeed, Pucha et al. (2019) have shown that its evolved stellar population extends to , while there is growing evidence that MW satellites have extended stellar halos (e.g. Chiti et al. 2021; Longeard et al. 2023; Sestito et al. 2023a, b; Jensen et al. 2024).
We continue our analysis by selecting sources with a , creating a high fidelity sample without the need to recalculate individual membership probabilities. This allows us to explore the kinematic properties of different sub-samples and make a detailed comparison with the HI kinematic field.
4.1 Kinematic properties along each pointing
| Sample | N | |||
| (km s) | (km s) | |||
| F1 | 410 | |||
| Main | F2 | 274 | ||
| F3 | 62 | |||
| F1 | 331 | |||
| RGS | F2 | 230 | ||
| F3 | 54 | |||
| F1 | 79 | |||
| MS | F2 | 44 | ||
| F3 | 8 | |||
Here we determine the kinematic properties of IC 1613’s stellar component as a function of radius by analysing the l.o.s. velocities of probable members within each MUSE pointing independently using a dispersion-only model. Results are reported in Table 4 and shown in Fig. 8, where we also compare them to the kinematic properties of the HI component, derived in Read et al. (2016) from the Little-THINGS survey data (Hunter et al. 2012; Oh et al. 2015). To make a direct comparison between gas and stars, we use the same central coordinates, in this case those of the HI, as there is a slight offset with the optical values (see Table 1)151515We note that throughout the text, velocities with the sub-script rot are the observed rotational velocities, neither corrected for the inclination nor for the asymmetric drift; those with the sub-script circ are circular velocities, thereby corrected for both effects.. We verified that the bias introduced by the rotation in the recovery of the velocity dispersion values in each field is negligible in our case (see also Leaman et al. 2012).
It is evident that the typical velocities of stars in each field change with respect to the systemic velocity when moving outward from the galaxy’s centre (see the left panels of Fig. 8, where we define the rotational velocity as as in order to show a positive trend). F1 and F2 contain almost 90% of the inspected stars and thus dominate the linear rotation signal recovered in the analysis of the entire sample (shown as a black solid line in figure). Indeed, the velocity difference between them is km s which, considering that the average distance between F1 and F2 is arcmin, is in agreement with the observed velocity gradient of km s arcmin.
From Fig. 8, we see that F2 and F3 show approximately the same rotational velocity (since their km s). Due to the limited spatial sampling, it is difficult to determine whether the velocity signal has reached a constant value or whether it will continue to vary radially. We recall that the only available spectroscopic dataset that we could use to improve the spatial sampling is that published by Kirby et al. (2014) which, unluckily, is mainly distributed along the optical minor axis of the galaxy providing, as already shown by Wheeler et al. (2017), a poor constraint to the rotation signal. A similar behaviour at these radii is also seen in the HI component, and it appears to be only a local flattening of the rotation curve, as it becomes clear when looking at the HI kinematics over a larger radial range (see also Oh et al. 2015; Read et al. 2016).
Inspection of the velocity dispersion of the stellar component reveals a radial decrease, with the of the central pointing being close to the value obtained from the full sample, while for F2 and F3 we find a significant decrease to a roughly similar lower value. The velocity dispersion profile of the HI starts at km s near the centre of the galaxy, then rises up to km s at (i.e., around the position of F2), to finally decrease and stabilise around km s for . The velocity dispersion of the HI is systematically lower than that obtained for F2 and F3 at the same radii, while this difference increases significantly when compared to the F1 value. It is probable that the lower central HI density due to the presence of a shell-like gas structure could have had an impact on the gas kinematics, and hence on the velocity dispersion profile, inside (see Moiseev & Lozinskaya 2012; Stilp et al. 2013). The HI velocity dispersion is indeed correlated to the gas surface mass density, which regulates the amount of local turbulence (e.g. Bacchini et al. 2020). We will investigate further in the future whether this may be related to the greater velocity dispersion measured from the stars in F1.
In order to estimate the galaxy’s circular velocity and directly compare the kinematics of stars and gas, it was necessary to remove the random motion component that suppresses the rotation curve. We followed the formalism of Read et al. (2016, using their Eqs. 9 and 12) to apply the so-called asymmetric drift correction. Although their analysis focused on the HI component, for our purposes the same reasoning could be applied to the stellar field. We also assumed the same inclination angle (see Table 1) for both tracers. After applying this correction, the stellar and HI components showed excellent agreement. This indicates that the differences observed in the rotational velocities and velocity dispersion between the HI and stellar components are indeed real.
We now examine the ratio of rotational to dispersion support for both stars and gas in IC 1613. For the stellar sample, we obtain a at the projected half-light radius , depending on whether we adopt as the average of the values between F2 and F3 or whether we extrapolate the values obtained for the linear rotation model reported in Table 3. We can consider the values from these two approaches as compatible with each other, taking into account the errors, as shown in Fig. 8. This is the observed without taking into account the inclination angle of IC 1613.
Using low-resolution HI data, Lake & Skillman (1989) reported a value of , in very good agreement with the best-fitting value from Read et al. (2016, see Table 1), which we have adopted. A similar value was also obtained from Oh et al. (2015), although unconstrained. However, Read et al. (2016) suggested that the true inclination could be as low as , allowing them for a better dynamical modelling. This is in contrast to Lake & Skillman (1989), who argue for an inclination angle of . This conclusion is based on the fact that the external isophotes of the HI agree well in shape and orientation with those of the stellar component, as is also the case between the spatial and kinematic P.A. of the HI. If the galaxy was seen almost face-on, it would be difficult to explain such features.
Depending on the inclination adopted, the stellar is between 0.8 and 1.1 for or between 1.9 and 2.7 for . The HI, for comparison, has an observed at the same radius, which rises to when the inclination angles are taken into account.
We can compare our results with those of two similarly bright LG dwarf irregular galaxies, WLM and NGC 6822. Unlike IC 1613, these are seen almost edge-on. WLM shows a stellar at , with the rotation axis aligned with that of the HI (Leaman et al. 2012). In contrast, the kinematics of NGC 6822 is more complex, as is the spatial distribution of its tracers. In fact, the P. A. of the HI is roughly perpendicular to the optical, and while both the stellar and HI components show signs of rotation (Weldrake et al. 2003; Demers et al. 2006; Swan et al. 2016; Wheeler et al. 2017; Belland et al. 2020), it is not clear whether they share the same kinematic axis. In any case, the stellar rotation signal seems to be between 0.4 and 1 (Swan et al. 2016; Wheeler et al. 2017; Belland et al. 2020). For both systems, the HI rotation resulted in much higher values with within for NGC 6822 (Weldrake et al. 2003), and for WLM (Oh et al. 2015; Iorio et al. 2017), similar to the IC 1613 case (assuming though the lowest inclination angle). Our results seem to be in good agreement with the conclusions drawn by Wheeler et al. (2017) that the gas-rich systems of the LG are not strong rotators in their stellar components (i.e. they have ).
4.2 Kinematic properties as a function of stellar type
We now divide our main sample of probable member stars into two sub-samples, young MS stars and evolved red giant stars, to examine the kinematic properties as a function of stellar type. Relying on the analysis in Sect. 3.1, we perform a simple selection according to the measured : red giant stars (RGS, i.e., belonging to the RGB and AGB with stellar ages Gyr) as those with K, and MS stars (having Gyr) as those with K. We verified that the limit is in general effective in separating the two samples, comparing with the expected from the colours of our targets using the empirical calibration for red giant stars from Alonso et al. (1999), valid in the colour range of and mostly metallicity independent.
The results are reported in Table 4 and Fig. 8. The results for the RGS are very similar (within 1-) to those of the full sample in terms of rotation velocities and velocity dispersion per field, as well as the global values of the linear velocity gradient. This is expected, as this sub-sample represents a large fraction of the main one. As is evident from Fig. 8, a model allowing for a flat rotation curve, or a more steeply rising rotation curve than the one we adopted, would have been a much better match to the kinematics of the MS stars. At all radii, the MS stars are showing a higher amplitude in their rotation with respect to the RGS (and the full sample), and in F2 and F3 also a significantly lower velocity dispersion. From the same figure, we can see that the MS stars follow the HI kinematic properties very closely (apart from the velocity dispersion in F1), while on the other hand, the kinematics of the older RGS, have decoupled from those of the HI. Therefore, the kinematics of the young MS stars appear to be still coupled to that of the HI component.
We checked for the presence of possible biases that could alter the recovery of the kinematic properties of the MS stars. On one hand, the region approximately separating the MS from the RGS, (see Fig. 4), contains the instability strip of IC 1613, where it is expected to find mainly classical Cepheids with short-term variability (i.e., with log and Gyr; Udalski et al. 2001; Bernard et al. 2010, see e.g.). Since we observe them in random phases, we expect any differences to be averaged out. It is also true that, especially for Cepheids, hydrogen or metallic lines can lead to different velocities (see, e.g., Vinko et al. 1998). Nonetheless, we find no significant differences when calculating the average velocities and velocity dispersion including or excluding stars in that color range. Similar considerations can be made for the Be-star candidates in our sample, which should also be short-period variables (e.g., Porter & Rivinius 2003; Rivinius et al. 2013), although in this case it is mostly the intensity of the emission lines that is affected. We repeated the same test finding values compatible within the errors. As for the RGS sample, this contains a small number of C-star candidates, which are long period variables (e.g., Menzies et al. 2015). However, we have verified that the C-stars have similar properties to the rest of the sample within the errors.
We did not check for the presence of binary stars, which could particularly affect the velocity dispersion values. The low S/N of the individual exposures for each field in our sample and the short period over which they were observed (between a few days and a month), preclude a quantitative analysis. We can, however, make qualitative estimates. For a sample of mostly RGB stars with high measured velocity dispersion ( km s), as it is our case, the impact of unresolved binaries is expected to be minimal (i.e. of inflation; Minor et al. 2010; Minor 2013; Spencer et al. 2017, 2018; Arroyo-Polonio et al. 2023, see). However, the case is different when focusing on the MS stars sample. The multiplicity fraction (i.e. the fraction of stars having at least one companion) is generally expected to be for A-types, which increase to for OB stars (e.g. Moe & Di Stefano 2017; Offner et al. 2023). In particular, these stars are often found in triples or higher order multiples. For Be stars such figures are less constrained, but mass accretion through binary interactions has been proposed as a possible channel for their formation (Pols et al. 1991; de Mink et al. 2013; Shao & Li 2014; Bodensteiner et al. 2020).
Considering that the multiple OBs cover a range of orbital periods, preferably a few days or a few decades (Moe & Di Stefano 2017), the stacking process to generate the spectra should already average the short-period ones. For the long-period ones (i.e. log) we expect velocity variations of km s, which could translate into a velocity dispersion inflation of (following Minor et al. 2010, assuming that they behave as short-period low-mass binaries and have intrinsic km s). This would especially affect the central pointing, where the fraction of OB stars is high (). It could thus partly explain why in this field the value of the velocity dispersion of the MS stars is comparable to that of the RGS, while in the other fields it tends to be closer to that of the HI. Another possible source of disturbance for the MS stars in F1 is that they may still be coupled to the motion of the ionised bubble, which is probably expanding (with a typical velocity of km s, Pokhrel et al. 2020). We also do not exclude the presence of possible runaway OB stars that could further impact the kinematics of the sample (e.g. Stoop et al. 2024).
In summary, the different examined tracers point to a similar picture: both IC 1613’s gaseous and stellar components are rotating, with the older stars decoupled from the HI motion mainly by the effect of asymmetric drift, while on timescales of Gyr, the stars have maintained a similar kinematics as the neutral gas. The velocity dispersion profile decreases with radius for all tracers, although in F1 the value of the MS stars may be affected by the presence of binaries or by the expanding motion of the central hole, which may have also affected the central HI velocity dispersion. Nevertheless, there seems to be a correlation between the velocity dispersion of the different tracers and their age, with values increasing as we move towards the older stars. This is in qualitative agreement with the results of Leaman et al. (2017), suggesting that stars are born from gas that becomes less turbulent over time due to decreasing gas fractions.
4.3 Dynamical mass estimation
We can use the results from the previous sections to estimate IC 1613’s dynamical mass, based on the stellar rotational velocities, corrected for the galaxy’s inclination, and the asymmetric drift; therefore turned into circular velocities. We perform this estimate within the 3D de-projected half-light radius to make a direct comparison with values from mass estimators for dispersion supported systems.
We took the average of the observed values of and between F2 and F3, assuming that both the rotation and dispersion profiles are rather flat up to the half-light radius. Assuming both values quoted in Read et al. (2016) for the the galaxy’s inclination i, we would have km s and km s, respectively. The dynamical mass is then between and , respectively. The corresponding mass-to-light ratio within the half-light radius would be between and .
For comparison, the Wolf et al. (2010) mass-estimator for dispersion-supported systems, that is , where G is the gravitational constant, yields within the same radius. This is within the previous results accounting for inclination, but also with the value reported by Kirby et al. (2014). We note that this mass estimator is valid as long as the system is spherically symmetric, without rotation and with a flat l.o.s. velocity dispersion profile. To some degree, all these conditions are violated for IC 1613.
It is important to note that the estimates of the dynamical mass, and dynamical mass-to-light ratio, based on rotational velocities given in this section should be considered as lower limits. The HI circular velocity curve of IC 1613 is known to rise very slowly, even with respect to the expectations of a cored dark matter density profile and assuming the more favourable inclination of . This appears to be the consequence of depressions, particularly evident around 0.5 and 1.5 kpc, due to the presence of large HI bubbles, as discussed in Read et al. (2016). If the stellar velocities reflect a similar behaviour, this would imply that also the stellar rotation curve suffers from these dips, thereby leading to an underestimate of the circular velocity and of the dynamical mass.
5 Metallicity properties
| Starkenburg et al. (2010) | Carrera et al. (2013) | ||||||||
| Sample | N | ||||||||
| (dex) | (dex) | (dex) | (dex ) | (dex) | (dex) | (dex) | (dex ) | ||
| Full | 275 | 0.29 | 0.26 | 0.37 | 0.34 | ||||
| F1 | 150 | 0.29 | 0.26 | 0.36 | 0.32 | ||||
| F2 | 93 | 0.31 | 0.26 | 0.35 | 0.36 | ||||
| F3 | 27 | 0.22 | 0.21 | 0.30 | 0.32 | ||||
We determined metallicities ([Fe/H]) for the individual stars from the equivalent widths of the near-IR Ca II triplet (CaT) lines. Since the available calibrations are developed for RGB stars, we will be restricting our analysis to this sub-sample (i.e., , with , which helps separating them from the hotter evolving giants). We verified the goodness of this selection using a set of isochrones (Bressan et al. 2012) with ([Fe/H] dex) and Gyr, which traced the lower range of metallicities and stellar ages for the RGB stars as obtained from the SFH analysis of IC 1613 (Skillman et al. 2014). We also selected all stars with S/N to keep the uncertainties on the equivalent widths (EWs) less than 20%. This resulted in a sub-sample of 275 input sources.
We made use of the Starkenburg et al. (2010) calibration. As shown in Kacharov et al. (2017), this calibration can be safely applied to data of spectral resolution as low as , comparable to that of our MUSE spectra around the CaT lines. The Starkenburg et al. calibration combines the EWs of the two reddest Ca II lines with the term, where V is the visual magnitude of the selected star and that of the horizontal branch. The EWs were obtained by fitting a Voigt profile over a 15 Å window around the selected Ca II lines, weighting each pixel value for the flux uncertainty stored in the error spectrum. The EW uncertainties were then calculated from the covariance matrix of the fitted Voigt parameters. We adopted a value of (Bernard et al. 2010). Final uncertainties for the metallicity values were obtained by error propagation of the EW uncertainties. We verified that the input magnitudes’ errors do not have a significant impact on the final [Fe/H] uncertainties.
The uncertainties obtained in the EW measurements resulted to be about 0.5 Å, while the average [Fe/H] error was dex. However, a few targets () have metallicity values with errors dex. We note that a small fraction of the stars in our sub-sample are likely to be AGB stars, brighter on average, for which the calibration method can be applied anyway, since it does not introduce a significant systematic bias compared to RGB stars (see discussion in Pont et al. 2004).
From the analysis of the metallicity values we obtained a median dex, a median absolute deviation dex, and an intrinsic scatter dex (following Eq. 8 of Kirby et al. 2013), as also reported in Table 5. For comparison, we also applied the Carrera et al. (2013) calibration using the fitting windows defined by Cenarro et al. (2001). The main difference with the Starkenburg et al. calibration is that this one is empirical, whereas the former also used synthetic spectra. Nevertheless, we obtained comparable results: median dex, dex and dex (see again Table 5). The recovered median metallicity is dex lower, but the offset is within the scatter of the data, in this case dex higher.
Our results are in good agreement with those reported by Kirby et al. (2013): median dex and dex, also obtained targeting the RGB stars of the galaxy. The small deviations we found, again of the order of 0.1 dex, can be attributed to the different spectral resolution between the data and implemented technique to obtain the metallicity values (namely by directly fitting the available Fe lines).
We further analysed our sample by considering each pointing separately. In Fig. 9, we show the metallicity distribution per field, for clarity only plotting values obtained with the Starkenburg et al. calibration. Visual inspection of the MDFs could give the impression of a decrease in the presence of metal-rich stars from F1 to F3 fields. Therefore, we ran a two-sample two-sided Kolmogorov-Smirnov test, comparing separately F2 and F3 against F1, which is the largest sample and the one with the widest range of metallicities. We found that it is likely that the samples are drawn from the same parent distribution (p-values higher than 0.05), therefore the lack of metal-rich stars in F2 and F3 is due to the lower statistics. This is consistent with IC 1613 constantly forming stars in a spatially homogeneous manner (see Skillman et al. 2014), at least out to , which is how far our data extend.
We found that the median [Fe/H] tends to decrease very slightly from the central to the outer field, as reported in Table 5 for both calibrations used (see also the right panel of Fig. 9). Thus, we investigated the presence of a radial metallicity gradient performing a Gaussian process regression (GPR) analysis using a Gaussian kernel together with a noise component to account for the intrinsic metallicity scatter of the data. Details of this type of analysis have been described extensively in Taibi et al. (2022). For both calibrations we obtained a value of the gradient that is consistent with zero within the uncertainties (see Fig. 9 and Table 5), meaning that the bulk of the data, with [Fe/H] values of dex, do not show significant spatial variation. This result is consistent with the conclusions of Taibi et al. (2022), who found that LG dwarf galaxies, except for those likely to have undergone past mergers, have similar gradient values.
The CaT method only applies to RGB stars, while we have mentioned that spexxy provides spectral parameters, including [Fe/H], for all stars in our sample. However, we have shown in the Appendix B.1 that it has limitations in providing correct estimates. Nevertheless, we were able to show qualitatively that the RGB stars should be at least 0.2 dex more metal-poor than the MS stars, as expected from the SFH of IC 1613 (Skillman et al. 2014) and spectroscopic measurements of young supergiants (Bresolin et al. 2007; Tautvaišienė et al. 2007; Berger et al. 2018).
6 Summary and future work
In this paper, we present results from the chemo-kinematic analysis of the stellar component of the isolated dwarf irregular galaxy IC 1613. The analysis is based on a new set of spectroscopic data obtained with 3 pointings of the VLT/MUSE instrument. We extracted sources, from which we separated the stellar objects for their subsequent spectral classification and analysis.
The quality of the dataset allowed for a generally accurate spectral classification (i.e., with determined to better than 500 K) of 808 stars. We found a majority of K-type stars, and representatives from all the other types (from O- to M-). The classification included the identification of a sample of 24 probable Be stars, together with a sample of 14 probable C stars.
For all types except the Be stars, the l.o.s. velocities have been determined with a pixel-by-pixel analysis, resulting in measurements with a mean precision of km and S/N For the Be stars we have determined the velocities by cross-correlation with dedicated templates. From this sample we identified 746 probable members (i.e., having ), which show a significant rotation pattern (when modelled as a linear velocity gradient, this yields km s arcmin). This is the first time that rotation for the stellar component is detected with high significance for this galaxy.
We further analysed the probable members looking at the kinematic properties of each individual field and for different stellar populations. The individual fields analysis confirmed the presence of a rotation signal, and the decrease with radius of the velocity dispersion values that appears to flatten moving outward. We found a general agreement with the velocity field of the neutral HI component, as analysed in Read et al. (2016), though, on average, the stellar values showed higher velocity dispersion and a slower rotation, which can be ascribed to asymmetric drift. The kinematic analysis of the different selected stellar populations showed a coherent picture with the analysis of the main sample, with the MS stars following more closely the kinematics of the HI gas than the evolved red giant stars. It appears that the kinematics of this younger stellar component is still coupled to that of the gas.
The chemical analysis was conducted on the sub-sample of RGB stars exploiting the wavelength region of the Ca II triplet lines. We obtained a median dex, with scatter values of dex and dex, in fairly good agreement with previous results reported by Kirby et al. (2013). The average [Fe/H] is well within the rms scatter of the stellar luminosity-metallicity relation found for the LG dwarf galaxies by the same authors. The spatial distribution of the [Fe/H] values showed no significant signs of a radial metallicity gradient, compatible with results from other LG dwarf galaxies (Leaman et al. 2013; Taibi et al. 2022).
In future dedicated efforts, we plan to perform a dedicated analysis of the massive stars in our sample using the FASTWIND stellar atmosphere code (Puls et al. 2005), which will allow us to obtain reliable stellar parameters from their spectra. We also aim to improve the identification process of the M-type stars in our sample, and therefore to constrain the spectroscopic C/M ratio. This will tell us more of the chemical properties of the intermediate-age stellar population of IC 1613. Finally, this MUSE data-set could also be used to analyse the ionized gas medium using available emission lines, like or nebular oxygen lines in order to get an insight on the kinematics of the shell-like gas structure visible in field F1 and compare it with that of the youngest stars.
Acknowledgements.
We thank M.-R. Cioni, L. R. Patrick and J. Read for useful discussions and comments. ST acknowledges funding of a Leibniz-Junior Research Group (PI: M. Pawlowski; project number J94/2020) via the Leibniz Competition. GB and CG acknowledge support from the Agencia Estatal de Investigación del Ministerio de Ciencia en Innovación (AEI-MICIN) and the European Regional Development Fund (ERDF) under Grant Number PID2020-118778GB-I00/10.13039/50110001103 and the AEI under Grant Number CEX2019-000920-S. GI acknowledges financial support from the European Research Council (ERC) through the ERC Consolidator Grant DEMOBLACK, under contract No. 770017. SK acknowledges funding from UKRI in the form of a Future Leaders Fellowship (grant no. MR/T022868/1) This research has made use of NASA’s Astrophysics Data System, VizieR catalogue access tool (CDS, Strasbourg, France, DOI: 10.26093/cds/vizier), and extensive use of Python3.8 (Van Rossum & Drake 2009), including iPython (v8.3, Pérez & Granger 2007), Numpy (v1.21, Harris et al. 2020), Scipy (v1.7, Virtanen et al. 2020), Matplotlib (v3.5, Hunter 2007), Astropy (v5.0, Astropy Collaboration et al. 2022) and Scikit-learn (v1.0, Pedregosa et al. 2011) packages.References
- Albert et al. (2000) Albert, L., Demers, S., & Kunkel, W. E. 2000, AJ, 119, 2780
- Alonso et al. (1999) Alonso, A., Arribas, S., & Martínez-Roger, C. 1999, A&AS, 140, 261
- Arroyo-Polonio et al. (2023) Arroyo-Polonio, J. M., Battaglia, G., Thomas, G. F., et al. 2023, A&A, 677, A95
- Asplund et al. (2009) Asplund, M., Grevesse, N., Sauval, A. J., & Scott, P. 2009, ARA&A, 47, 481
- Astropy Collaboration et al. (2022) Astropy Collaboration, Price-Whelan, A. M., Lim, P. L., et al. 2022, ApJ, 935, 167
- Bacchini et al. (2020) Bacchini, C., Fraternali, F., Iorio, G., et al. 2020, A&A, 641, A70
- Bacon et al. (2014) Bacon, R., Vernet, J., Borisova, E., et al. 2014, The Messenger, 157, 13
- Battaglia & Nipoti (2022) Battaglia, G. & Nipoti, C. 2022, Nature Astronomy, 6, 659
- Battaglia et al. (2022) Battaglia, G., Taibi, S., Thomas, G. F., & Fritz, T. K. 2022, A&A, 657, A54
- Belland et al. (2020) Belland, B., Kirby, E., Boylan-Kolchin, M., & Wheeler, C. 2020, ApJ, 903, 10
- Bennet et al. (2023) Bennet, P., Patel, E., Sohn, S. T., et al. 2023, arXiv e-prints, arXiv:2312.09276
- Berger et al. (2018) Berger, T. A., Kudritzki, R.-P., Urbaneja, M. A., et al. 2018, ApJ, 860, 130
- Bernard et al. (2007) Bernard, E. J., Aparicio, A., Gallart, C., Padilla-Torres, C. P., & Panniello, M. 2007, AJ, 134, 1124
- Bernard et al. (2010) Bernard, E. J., Monelli, M., Gallart, C., et al. 2010, ApJ, 712, 1259
- Bodensteiner et al. (2020) Bodensteiner, J., Shenar, T., & Sana, H. 2020, A&A, 641, A42
- Borissova et al. (2000) Borissova, J., Georgiev, L., Kurtev, R., et al. 2000, Rev. Mexicana Astron. Astrofis., 36, 151
- Borissova et al. (2004) Borissova, J., Kurtev, R., Georgiev, L., & Rosado, M. 2004, A&A, 413, 889
- Bresolin et al. (2007) Bresolin, F., Urbaneja, M. A., Gieren, W., Pietrzyński, G., & Kudritzki, R.-P. 2007, ApJ, 671, 2028
- Bressan et al. (2012) Bressan, A., Marigo, P., Girardi, L., et al. 2012, MNRAS, 427, 127
- Buchner et al. (2014) Buchner, J., Georgakakis, A., Nandra, K., et al. 2014, A&A, 564, A125
- Buck et al. (2019) Buck, T., Macciò, A. V., Dutton, A. A., Obreja, A., & Frings, J. 2019, MNRAS, 483, 1314
- Bundy et al. (2015) Bundy, K., Bershady, M. A., Law, D. R., et al. 2015, ApJ, 798, 7
- Carrera et al. (2013) Carrera, R., Pancino, E., Gallart, C., & del Pino, A. 2013, MNRAS, 434, 1681
- Cenarro et al. (2001) Cenarro, A. J., Cardiel, N., Gorgas, J., et al. 2001, MNRAS, 326, 959
- Chambers et al. (2016) Chambers, K. C., Magnier, E. A., Metcalfe, N., et al. 2016, arXiv e-prints, arXiv:1612.05560
- Chiti et al. (2021) Chiti, A., Frebel, A., Simon, J. D., et al. 2021, Nature Astronomy, 5, 392
- Chun et al. (2015) Chun, S.-H., Jung, M., Kang, M., Kim, J.-W., & Sohn, Y.-J. 2015, A&A, 578, A51
- Chun et al. (2022) Chun, S.-H., Yoon, S.-C., Oh, H., Park, B.-G., & Hwang, N. 2022, ApJ, 939, 28
- Cole et al. (1999) Cole, A. A., Tolstoy, E., Gallagher, John S., I., et al. 1999, AJ, 118, 1657
- Collins & Read (2022) Collins, M. L. M. & Read, J. I. 2022, Nature Astronomy, 6, 647
- Cortese et al. (2014) Cortese, L., Fogarty, L. M. R., Ho, I. T., et al. 2014, ApJ, 795, L37
- de Mink et al. (2013) de Mink, S. E., Langer, N., Izzard, R. G., Sana, H., & de Koter, A. 2013, ApJ, 764, 166
- Demers et al. (2006) Demers, S., Battinelli, P., & Kunkel, W. E. 2006, ApJ, 636, L85
- Feroz et al. (2009) Feroz, F., Hobson, M. P., & Bridges, M. 2009, MNRAS, 398, 1601
- Gaia Collaboration et al. (2023) Gaia Collaboration, Vallenari, A., Brown, A. G. A., et al. 2023, A&A, 674, A1
- Gallart et al. (1996) Gallart, C., Aparicio, A., & Vilchez, J. M. 1996, AJ, 112, 1928
- Gallart et al. (2015) Gallart, C., Monelli, M., Mayer, L., et al. 2015, ApJ, 811, L18
- Garcia et al. (2014) Garcia, M., Herrero, A., Najarro, F., Lennon, D. J., & Alejandro Urbaneja, M. 2014, ApJ, 788, 64
- Garcia et al. (2009) Garcia, M., Herrero, A., Vicente, B., et al. 2009, A&A, 502, 1015
- Georgiev et al. (1999) Georgiev, L., Borissova, J., Rosado, M., et al. 1999, A&AS, 134, 21
- Girardi et al. (2000) Girardi, L., Bressan, A., Bertelli, G., & Chiosi, C. 2000, A&AS, 141, 371
- Gonneau et al. (2016) Gonneau, A., Lançon, A., Trager, S. C., et al. 2016, A&A, 589, A36
- Gull et al. (2022) Gull, M., Weisz, D. R., Senchyna, P., et al. 2022, ApJ, 941, 206
- Hamren et al. (2016) Hamren, K., Beaton, R. L., Guhathakurta, P., et al. 2016, ApJ, 828, 15
- Harris et al. (2020) Harris, C. R., Millman, K. J., van der Walt, S. J., et al. 2020, Nature, 585, 357–362
- Higgs et al. (2021) Higgs, C. R., McConnachie, A. W., Annau, N., et al. 2021, MNRAS, 503, 176
- Holtzman et al. (2006) Holtzman, J. A., Afonso, C., & Dolphin, A. 2006, ApJS, 166, 534
- Hunter et al. (2012) Hunter, D. A., Ficut-Vicas, D., Ashley, T., et al. 2012, AJ, 144, 134
- Hunter (2007) Hunter, J. D. 2007, Computing in science & engineering, 9, 90
- Husser (2012) Husser, T.-O. 2012, PhD thesis, Georg August University of Gottingen, Germany
- Husser et al. (2013) Husser, T. O., Wende-von Berg, S., Dreizler, S., et al. 2013, A&A, 553, A6
- Iorio et al. (2017) Iorio, G., Fraternali, F., Nipoti, C., et al. 2017, MNRAS, 466, 4159
- Jensen et al. (2024) Jensen, J., Hayes, C. R., Sestito, F., et al. 2024, MNRAS, 527, 4209
- Júlio et al. (2023) Júlio, M. P., Brinchmann, J., Zoutendijk, S. L., et al. 2023, A&A, 678, A38
- Kacharov et al. (2017) Kacharov, N., Battaglia, G., Rejkuba, M., et al. 2017, MNRAS, 466, 2006
- Kamann et al. (2013) Kamann, S., Wisotzki, L., & Roth, M. M. 2013, A&A, 549, A71
- Kirby et al. (2014) Kirby, E. N., Bullock, J. S., Boylan-Kolchin, M., Kaplinghat, M., & Cohen, J. G. 2014, MNRAS, 439, 1015
- Kirby et al. (2013) Kirby, E. N., Cohen, J. G., Guhathakurta, P., et al. 2013, ApJ, 779, 102
- Koleva et al. (2009) Koleva, M., Prugniel, P., Bouchard, A., & Wu, Y. 2009, A&A, 501, 1269
- Kroupa (2001) Kroupa, P. 2001, MNRAS, 322, 231
- Lake & Skillman (1989) Lake, G. & Skillman, E. D. 1989, AJ, 98, 1274
- Leaman et al. (2017) Leaman, R., Mendel, J. T., Wisnioski, E., et al. 2017, MNRAS, 472, 1879
- Leaman et al. (2013) Leaman, R., Venn, K. A., Brooks, A. M., et al. 2013, ApJ, 767, 131
- Leaman et al. (2012) Leaman, R., Venn, K. A., Brooks, A. M., et al. 2012, ApJ, 750, 33
- Lee et al. (2003) Lee, H., Grebel, E. K., & Hodge, P. W. 2003, A&A, 401, 141
- Longeard et al. (2023) Longeard, N., Jablonka, P., Battaglia, G., et al. 2023, MNRAS, 525, 3086
- Lozinskaya et al. (2003) Lozinskaya, T. A., Moiseev, A. V., & Podorvanyuk, N. Y. 2003, Astronomy Letters, 29, 77
- McConnachie (2012) McConnachie, A. W. 2012, AJ, 144, 4
- McConnachie et al. (2021) McConnachie, A. W., Higgs, C. R., Thomas, G. F., et al. 2021, MNRAS, 501, 2363
- McLeod et al. (2020) McLeod, A. F., Kruijssen, J. M. D., Weisz, D. R., et al. 2020, ApJ, 891, 25
- McQuinn et al. (2017) McQuinn, K. B. W., Boyer, M. L., Mitchell, M. B., et al. 2017, ApJ, 834, 78
- Menzies et al. (2015) Menzies, J. W., Whitelock, P. A., & Feast, M. W. 2015, MNRAS, 452, 910
- Minor (2013) Minor, Q. E. 2013, ApJ, 779, 116
- Minor et al. (2010) Minor, Q. E., Martinez, G., Bullock, J., Kaplinghat, M., & Trainor, R. 2010, ApJ, 721, 1142
- Moe & Di Stefano (2017) Moe, M. & Di Stefano, R. 2017, ApJS, 230, 15
- Moiseev & Lozinskaya (2012) Moiseev, A. V. & Lozinskaya, T. A. 2012, MNRAS, 423, 1831
- Offner et al. (2023) Offner, S. S. R., Moe, M., Kratter, K. M., et al. 2023, in Astronomical Society of the Pacific Conference Series, Vol. 534, Protostars and Planets VII, ed. S. Inutsuka, Y. Aikawa, T. Muto, K. Tomida, & M. Tamura, 275
- Oh et al. (2015) Oh, S.-H., Hunter, D. A., Brinks, E., et al. 2015, AJ, 149, 180
- Pedregosa et al. (2011) Pedregosa, F., Varoquaux, G., Gramfort, A., et al. 2011, Journal of Machine Learning Research, 12, 2825
- Pérez & Granger (2007) Pérez, F. & Granger, B. E. 2007, Computing in Science & Engineering, 9
- Pokhrel et al. (2020) Pokhrel, N. R., Simpson, C. E., & Bagetakos, I. 2020, AJ, 160, 66
- Pols et al. (1991) Pols, O. R., Cote, J., Waters, L. B. F. M., & Heise, J. 1991, A&A, 241, 419
- Pont et al. (2004) Pont, F., Zinn, R., Gallart, C., Hardy, E., & Winnick, R. 2004, AJ, 127, 840
- Porter & Rivinius (2003) Porter, J. M. & Rivinius, T. 2003, PASP, 115, 1153
- Pucha et al. (2019) Pucha, R., Carlin, J. L., Willman, B., et al. 2019, ApJ, 880, 104
- Puls et al. (2005) Puls, J., Urbaneja, M. A., Venero, R., et al. 2005, A&A, 435, 669
- Putman et al. (2021) Putman, M. E., Zheng, Y., Price-Whelan, A. M., et al. 2021, ApJ, 913, 53
- Read et al. (2016) Read, J. I., Iorio, G., Agertz, O., & Fraternali, F. 2016, MNRAS, 462, 3628
- Ren et al. (2022) Ren, T., Jiang, B., Ren, Y., & Yang, M. 2022, Universe, 8, 465
- Richard et al. (2019) Richard, J., Bacon, R., Blaizot, J., et al. 2019, arXiv e-prints, arXiv:1906.01657
- Rivinius et al. (2013) Rivinius, T., Carciofi, A. C., & Martayan, C. 2013, A&A Rev., 21, 69
- Robin et al. (2003) Robin, A. C., Reylé, C., Derrière, S., & Picaud, S. 2003, A&A, 409, 523
- Roth et al. (2018) Roth, M. M., Sandin, C., Kamann, S., et al. 2018, A&A, 618, A3
- Sánchez et al. (2012) Sánchez, S. F., Rosales-Ortega, F. F., Marino, R. A., et al. 2012, A&A, 546, A2
- Sánchez-Blázquez et al. (2006) Sánchez-Blázquez, P., Peletier, R. F., Jiménez-Vicente, J., et al. 2006, MNRAS, 371, 703
- Sandage (1971) Sandage, A. 1971, ApJ, 166, 13
- Schlafly & Finkbeiner (2011) Schlafly, E. F. & Finkbeiner, D. P. 2011, ApJ, 737, 103
- Schootemeijer et al. (2022) Schootemeijer, A., Lennon, D. J., Garcia, M., et al. 2022, A&A, 667, A100
- Sestito et al. (2023a) Sestito, F., Roediger, J., Navarro, J. F., et al. 2023a, MNRAS, 523, 123
- Sestito et al. (2023b) Sestito, F., Zaremba, D., Venn, K. A., et al. 2023b, MNRAS, 525, 2875
- Shao & Li (2014) Shao, Y. & Li, X.-D. 2014, ApJ, 796, 37
- Sibbons et al. (2015) Sibbons, L. F., Ryan, S. G., Irwin, M., & Napiwotzki, R. 2015, A&A, 573, A84
- Silich et al. (2006) Silich, S., Lozinskaya, T., Moiseev, A., et al. 2006, A&A, 448, 123
- Simon (2019) Simon, J. D. 2019, ARA&A, 57, 375
- Skillman et al. (2014) Skillman, E. D., Hidalgo, S. L., Weisz, D. R., et al. 2014, ApJ, 786, 44
- Skillman et al. (2003) Skillman, E. D., Tolstoy, E., Cole, A. A., et al. 2003, ApJ, 596, 253
- Spencer et al. (2018) Spencer, M. E., Mateo, M., Olszewski, E. W., et al. 2018, AJ, 156, 257
- Spencer et al. (2017) Spencer, M. E., Mateo, M., Walker, M. G., et al. 2017, AJ, 153, 254
- Starkenburg et al. (2010) Starkenburg, E., Hill, V., Tolstoy, E., et al. 2010, A&A, 513, A34
- Stetson (1987) Stetson, P. B. 1987, PASP, 99, 191
- Stilp et al. (2013) Stilp, A. M., Dalcanton, J. J., Skillman, E., et al. 2013, ApJ, 773, 88
- Stoop et al. (2024) Stoop, M., Derkink, A., Kaper, L., et al. 2024, A&A, 681, A21
- Swan et al. (2016) Swan, J., Cole, A. A., Tolstoy, E., & Irwin, M. J. 2016, MNRAS, 456, 4315
- Taibi et al. (2022) Taibi, S., Battaglia, G., Leaman, R., et al. 2022, A&A, 665, A92
- Taibi et al. (2020) Taibi, S., Battaglia, G., Rejkuba, M., et al. 2020, A&A, 635, A152
- Tautvaišienė et al. (2007) Tautvaišienė, G., Geisler, D., Wallerstein, G., et al. 2007, AJ, 134, 2318
- Tolstoy et al. (2009) Tolstoy, E., Hill, V., & Tosi, M. 2009, ARA&A, 47, 371
- Udalski et al. (2001) Udalski, A., Wyrzykowski, L., Pietrzynski, G., et al. 2001, Acta Astron., 51, 221
- Valdes et al. (2004) Valdes, F., Gupta, R., Rose, J. A., Singh, H. P., & Bell, D. J. 2004, ApJS, 152, 251
- van de Ven et al. (2006) van de Ven, G., van den Bosch, R. C. E., Verolme, E. K., & de Zeeuw, P. T. 2006, A&A, 445, 513
- van den Bergh (2000) van den Bergh, S. 2000, The Galaxies of the Local Group
- Van Rossum & Drake (2009) Van Rossum, G. & Drake, F. L. 2009, Python 3 Reference Manual (Scotts Valley, CA: CreateSpace)
- Vaz et al. (2023) Vaz, D., Brinchmann, J., Zoutendijk, S. L., et al. 2023, A&A, 678, A59
- Vazdekis et al. (2016) Vazdekis, A., Koleva, M., Ricciardelli, E., Röck, B., & Falcón-Barroso, J. 2016, MNRAS, 463, 3409
- Vazdekis et al. (2012) Vazdekis, A., Ricciardelli, E., Cenarro, A. J., et al. 2012, MNRAS, 424, 157
- Vinko et al. (1998) Vinko, J., Remage Evans, N., Kiss, L. L., & Szabados, L. 1998, MNRAS, 296, 824
- Virtanen et al. (2020) Virtanen, P., Gommers, R., Oliphant, T. E., et al. 2020, Nature Methods, 17, 261
- Walker et al. (2009) Walker, M. G., Mateo, M., Olszewski, E. W., Sen, B., & Woodroofe, M. 2009, AJ, 137, 3109
- Weilbacher et al. (2012) Weilbacher, P. M., Streicher, O., Urrutia, T., et al. 2012, Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 8451, Design and capabilities of the MUSE data reduction software and pipeline, 84510B
- Weldrake et al. (2003) Weldrake, D. T. F., de Blok, W. J. G., & Walter, F. 2003, MNRAS, 340, 12
- Wheeler et al. (2017) Wheeler, C., Pace, A. B., Bullock, J. S., et al. 2017, MNRAS, 465, 2420
- Wolf et al. (2010) Wolf, J., Martinez, G. D., Bullock, J. S., et al. 2010, MNRAS, 406, 1220
- Wolf (1906) Wolf, M. 1906, MNRAS, 67, 91
- Zoutendijk et al. (2020) Zoutendijk, S. L., Brinchmann, J., Boogaard, L. A., et al. 2020, A&A, 635, A107
Appendix A Calibration of MUSE catalogues
The raw catalogues obtained from the MUSE VRI pseudo-images contain the spatial and photometric information of all extracted sources, which is needed for the kinematic and chemical analysis we conducted. However, they needed astrometric and photometric calibration. For this purpose, we used the publicly available HST/WFPC2 catalogues from Holtzman et al. (2006), which corresponded to the F1 and F3 fields, and a proprietary Subaru/SuprimeCam catalogue for the F2 field, kindly provided to us by M. Monelli (IAC).
We note that the field-of-view (FoV) of the Subaru catalogue () is wide enough to cover all the three MUSE fields at once. However, it lacked an astrometric solution at first, and it showed severe crowding toward the galaxy’s centre, where F1 is placed. The HST catalogues, on the other hand, did not show any of these problems and we used them as a baseline during the astrometric and photometric calibration processes.
We downloaded the PSF-extracted HST catalogues with transformed magnitudes in the Johnson’s system171717Catalogues can be found at http://astronomy.nmsu.edu/holtz/archival/ic1613/html/ic1613.html. The covered photometric bands were V and I, with the addition of the B-band for the central HST pointing. The Subaru catalogue covered instead the Johnson’s B and V bands. These are deep photometric catalogues, reaching down to the horizontal branch (at ), which means that they cover a much wider photometric range than that of our MUSE data. Before proceeding, we performed a pre-cleaning, selecting all sources marked as stellar in the HST catalogues, while retaining those targets in the Subaru catalogue having DAOPHOT parameters sharpness (SHARP) between and , and a goodness-of-fit (CHI) .
The astrometric solution and photometric calibration were obtained using the suite of codes CataXcorr and CataComb, kindly provided to us by P. Montegriffo and M. Bellazzini (INAF-OAS). We started with the astrometric registration of the Subaru catalogue. Due to its large FoV compared to that of the HST catalogues, we first performed a pre-astrometric registration using the stars in common with a PanSTARRS catalogue (Chambers et al. 2016) generated within the FoV of the Subaru catalogue. We note that the PanSTARRS catalogue was shallower that the Subaru one, reaching only a magnitude lower the tip of the RGB (at ). We then used the external HST field catalogue to refine the astrometric solution, improving its accuracy to better than 0.05″.
The photometric bands of the MUSE catalogues we were interested in calibrating were the V and I bands. We recall that the raw catalogues extracted from the data cubes were made in the VRI bands. The B-band was not extracted because it is not completely covered by the wavelength range of MUSE. Since the HST catalogues covered the BVI and VI bands (central and outer pointing, respectively), while the Subaru only covered the BV bands, as a compromise, we decided to add the I-band to the Subaru catalogue and calibrate the MUSE catalogues in VI.
We inferred the Subaru I-band using the common targets with the central HST catalogue (which covered the B-band). We applied the following linear equation: . We obtained a colour term and a zero-point , which allowed to calculate the magnitudes, whose agreement with the of the common targets was good, showing a median absolute deviation (MAD) scatter of 0.08 mags. However, comparing the colours of the common sources, we found a difference of mags driven by the bluer objects. This implied the introduction of a colour bias for the probable MS stars of the F2 field, as can be appreciated in Fig. 3, whose calibration was based on the Subaru catalogue. However, their calibrated resulted safely below 0.5 mags (i.e. the colour limit of the RGB).
Finally, we used the mentioned catalogues to astrometrically and photometrically calibrate the corresponding MUSE fields – the HST ones with the F1 and F3 fields, while the Subaru one with the F2 field. For the photometric calibration of the MUSE pointings using the reference catalogues, we applied the following equations: , and . The accuracy of the astrometric and photometric solutions resulted on average 0.08″and 0.05 mags, respectively.
Appendix B Consistency checks
B.1 Verify SPEXXY general performance
We verified the performance of spexxy in recovering l.o.s. velocities and associated errors. We recall that spexxy performs a full spectral fit to the observed spectra using interpolated templates generated from the PHOENIX library of high-resolution synthetic spectra (Husser et al. 2013). It requires a configuration file specifying initial values for velocity, effective temperature, surface gravity, metallicity and [/M] from which to start the fit. We obtained the initial values from a set of isochrones (Girardi et al. 2000; Bressan et al. 2012) with dex (around the mean metallicity of IC 1613) by performing a linear regression between the observed colours and magnitudes and the theoretical and . For the spectral fitting, we used the available PHOENIX templates, convolved with the MUSE line spread function, with solar scale chemical composition, thus fixing dex. We let the code find the best fit of all other parameters (i.e., , and , together with the l.o.s. velocity). We also masked certain spectral regions to avoid possible emission lines due to the interstellar medium and residuals due to sky line subtraction (such as the persistent residual due to the [OI] line at 5577 Å). To test for possible systematic velocity shifts we ran the code on several noiseless templates from the PHOENIX library of giant stars at different (from 2500 K to 15000 K) at fixed dex. Templates were shifted at several l.o.s. velocities, from km s to km s at step of 50 km s. As initial guess for the velocity fitting, we set km s (i.e. approximately IC 1613 systemic velocity). We found that in general spexxy correctly recovers the assigned l.o.s. velocities (mean and scatter of the difference between input and output velocities being km s), although we noticed some failure around the extremes (i.e. for km s and km s) with velocities variations of up to 50%. This occurred in only 3% of the cases that are those differing the most from the initial guess for the velocity. We recall that based on the Besançon model we expect almost no contaminants with true velocities km s, while the variation at km s is small enough to keep such possible contaminants away from the bulk velocity of IC 1613.
Regarding the other parameters, values were also well recovered, except for some scatter of K at temperatures K; was recovered with a significant scatter ( dex) for the hottest stars and those with K (the latter a minority in our main sample); was also well recovered, but with a large scatter ( dex) for templates with extreme velocities. Again, this is mostly due to the large difference between the initial guess and the true velocity of the stellar template, which leads spexxy to an incorrect parameter estimate. Therefore, spexxy generally recovers accurate velocities and spectral parameters.
To test whether spexxy provides the correct uncertainties, we developed a series of Monte Carlo tests in which we generated mock spectra at different S/N, starting from noiseless templates of different spectral types and injecting noise directly from the error spectra associated to the observed ones. However, we found that the flux uncertainty associated with the observed spectra was underestimated by an average factor of 1.4 with respect to the measured noise on the observed flux. This was found systematically for all spectra, regardless of their spectral type or S/N. It is most likely a result of the resampling into a regularly sampled data cube during data reduction. This introduces covariances between pixels that are neglected by the pipeline. In principle, this factor is not a problem, as the uncertainties of spexxy are calculated from the residuals between the input spectrum and the best-fit template, but we had to take this into account when generating the mock spectra.
Since we have two main stellar populations in our observed sample, RGS and MS stars, we generated mock spectra representative of these two sub-samples. For the RGS, we selected stars with K, with a difference between observed values found by spexxy and ULySS of no more than 500 K. From their best-fitting templates and scaled noise spectra, we generated 250 mock spectra in bins of S/N around . The same was done for the MS stars, but for stars with K, generating mock spectra at S/N around . Results are reported in Table 6.
We found that for the RGS stars their velocities and associated errors are well recovered, with deviations between the mean of the errors and the measurement scatter within km s. The effective temperature was also well recovered, although the errors were as much as K lower than the measurement scatter. On the other hand, and were relatively more uncertain, with errors underestimated between 0.1 and 0.2 dex (higher deviations at lower S/N), leading to an offset on the recovered values of up to 0.1 dex.
For the MS stars we recovered higher deviations from the input values than the RGS case. In particular, velocity errors at S/N were underestimated by more than km s, while at the highest S/N the deviation was less than 1 km s. In the main text we reported that we made a quality cut in our sample by excluding stars with observed velocity errors km s. This cut mainly excluded MS stars with the lowest S/N, while ensuring that we kept those with well-estimated velocity errors within 10%. For the spectral parameters, the associated errors were generally more underestimated () than in the RGS case.
Therefore, it can be concluded that spexxy is generally effective in accurately and precisely recovering stellar velocities. However, it struggles to correctly recover stellar parameters, with being the most uncertain parameter in our mock tests. This issue is primarily related to the low resolution of the MUSE data rather than a problem with the fitting routine. It is also important to note that our tests only explored a limited region of the parameter space. To further evaluate the performance of spexxy in recovering values, we conducted an additional test by comparing the values obtained from spexxy with those obtained for the RGB stars using the CaT method (see details in Sect. 5).
We found a significant deviation between the distributions of values from the two methods of dex calculated between the medians, with those obtained with spexxy being metal richer. The recovered deviation is significantly larger than expected based on the mock tests (see again Table 6) and did not show a dependence with S/N. This discrepancy can be partly attributed to the poor recovery of log(g) by spexxy, as shown by the mock tests. In fact, we re-run spexxy on our targets while fixing log(g) to the initial input values, thus reducing the number of free parameters to fit. As a result, the offset to the CaT-based values decreased to dex. The observed difference remains significant, but the distributions now have comparable standard deviations.
Fixing the value of log(g) during the spexxy fitting did not significantly change the distribution of the MS stars. This could be due to the lower sensitivity of the MS stars to log(g) variations when recovering metallicity values. The results of the mock tests show that the recovered values are quite accurate (for S/N), but with slightly underestimated errors.
Assuming that fixing the log(g) yields metallicity values accurate to within a constant shift in the RGB case, we can make a relative comparison between the distribution of RGB and MS stars using spexxy outputs. We further refined the selection of MS stars by keeping the probable kinematic members with K (i.e. up to the grid limit of the PHOENIX library) and S/N to ensure a similar sample quality as for the RGB. This reduced the MS sample to 30 stars. Their average metallicity was comparable to that of the RGB sample, but with a larger associated standard deviation. Taking into account that the average value for the RGB may be overestimated by 0.2 dex, we can conclude that the MS stars are expected to be more metal-rich than the RGB. This difference may increase if we take into account the for the two populations, which we set to zero during the spexxy fitting. The fact that the MS stars are more metal-rich than the RGB would be expected (Skillman et al. 2014), since also spectroscopic measurements of young blue and red supergiants from the literature suggest an average dex (Bresolin et al. 2007; Tautvaišienė et al. 2007; Garcia et al. 2014; Berger et al. 2018), while the ionised HII medium may be slightly metal-rich (by dex, Bresolin et al. 2007, e.g.).
| S/N | ||||||||
| (pxl) | (km s) | (K) | (dex) | (dex) | ||||
| 3.5 | ||||||||
| 5 | ||||||||
| 10 | ||||||||
| 15 | ||||||||
| 25 | ||||||||
| 5 | ||||||||
| 10 | ||||||||
| 20 | ||||||||
| 50 |
B.2 Comparison between SPEXXY and ULySS
In our analysis, we used spexxy and ULySS in a complementary way: the former for l.o.s. velocity measurements, while the latter for spectral classification. However, since both codes provide similar outputs, we compared their performances in recovering the l.o.s. velocity and from the observed spectra. We considered all targets in the sample used for the kinematic analysis (see Sect. 4). The l.o.s. velocity errors for the ULySS measurements were obtained from Monte Carlo simulations using several templates of giant stars at different intervals of temperature and S/N. We made this choice because the formal errors returned by ULySS were generally overestimated, without following any clear trend with the S/N or spectral type. For the Monte Carlo simulations we chose seven templates from the MIUSCAT library (see Sect.3.1), with ranging from 3500 K to 20000 K. The simulations were performed by adding Poisson noise to each template to obtain simulated spectra at increasing S/N intervals, with values ranging from 4 to 30. We repeated each process 100 times. For each template, the recovered velocity scatter as a function of S/N was fitted by an exponential law. The ULySS measurements were then divided into effective temperature bins according to the templates considered, and the velocity errors assigned using the corresponding exponential fit. First we compared the velocity differences for pair measurements obtained with ULySS and spexxy. We found that % of the measured velocities are in agreement within the combined uncertainties, obtained as . The comparison is shown in the left panel of Fig. 10 where for clarity we reported only those measurements with a combined error km s. Nevertheless, the difference between the ULySS and spexxy measurements showed an average positive shift of km s. What is more, a higher shift of km s is observed for those objects with a S/N. This was somehow expected for the UlySS measurements, since the Monte Carlo tests showed average velocity shifts of up to 20 km s for templates with effective temperatures of K and greater than K having simulated . On the other hand, we have already showed in the previous section the capacity of spexxy to correctly recover l.o.s. velocities, even at the lowest S/N for red giant stars. This is a very important point because most of the stars in our dataset are RGS in a low S/N regime (), where spexxy showed its strength in the recovery of l.o.s. velocities. The comparison of effective temperatures, as shown in the right-hand panel of Fig. 10, gave good results for K, although the spexxy values tend to be systematically K higher than the ULySS values. At higher , spexxy values started to deviate significantly. This is mainly due to the fact that spexxy measurements are based on the PHOENIX library, which is not optimized for very hot stars. In the ULySS case, instead, for each spectrum the stored temperature value is that of the best fitting template, which was chosen by eye among the most likely templates. We recall that this procedure had an accuracy of up to K, for a quality PCL-flag between 3 and 4.
Appendix C Observing log and examples of extracted spectra
We report here a table with the observing log of our VLT/MUSE observations of IC 1613. We show as well some examples of the extracted spectra analysed in this work.
| Field | Position (RA, Dec) | Date / Hour | Exp. | Airmass | DIMM Seeing |
| (J2000) | (UT) | (s) | (arcsec) | ||
| F1 | 01:04:49.00, +02:07:16.3 | 2016-08-12 / 09:01 | 1385 | 1.14 | 0.79 |
| 01:04:49.03, +02:07:16.0 | 2016-08-12 / 09:26 | 1385 | 1.17 | 0.68 | |
| 01:04:49.00, +02:07:15.6 | 2016-08-13 / 07:57 | 1385 | 1.12 | 0.65 | |
| 01:04:48.97, +02:07:15.8 | 2016-08-13 / 08:23 | 1385 | 1.12 | 0.64 | |
| 01:04:49.03, +02:07:15.6 | 2016-08-14 / 08:29 | 1385 | 1.12 | 0.67 | |
| 01:04:49.00, +02:07:15.9 | 2016-08-14 / 08:54 | 1385 | 1.14 | 0.68 | |
| 01:04:48.97, +02:07:16.3 | 2016-09-02 / 03:40 | 1385 | 1.76 | 0.60 | |
| 01:04:48.98, +02:07:15.3 | 2016-09-02 / 04:05 | 1385 | 1.56 | 0.62 | |
| F2 | 01:04:40.69, +02:04:45.2 | 2016-09-28 / 04:00 | 1385 | 1.19 | 0.73 |
| 01:04:40.70, +02:04:46.1 | 2016-09-28 / 04:25 | 1385 | 1.15 | 0.54 | |
| 01:04:40.67, +02:04:45.6 | 2016-09-29 / 05:40 | 1385 | 1.13 | 0.52 | |
| 01:04:40.73, +02:04:45.9 | 2016-09-29 / 06:05 | 1385 | 1.15 | 0.56 | |
| 01:04:40.73, +02:04:45.4 | 2016-09-30 / 03:56 | 1385 | 1.18 | 0.69 | |
| 01:04:40.70, +02:04:45.7 | 2016-09-30 / 04:21 | 1385 | 1.14 | 0.71 | |
| 01:04:40.72, +02:04:45.2 | 2016-09-30 / 04:52 | 1385 | 1.12 | 0.76 | |
| 01:04:40.68, +02:04:45.9 | 2016-09-30 / 05:17 | 1385 | 1.12 | 0.73 | |
| F3 | 01:04:29.44, +02:03:35.3 | 2016-10-04 / 04:41 | 1385 | 1.12 | 0.67 |
| 01:04:29.39, +02:03:35.4 | 2016-10-04 / 05:06 | 1385 | 1.12 | 0.55 | |
| 01:04:29.42, +02:03:35.5 | 2016-10-04 / 05:36 | 1385 | 1.14 | 0.70 | |
| 01:04:29.42, +02:03:36.0 | 2016-10-04 / 06:01 | 1385 | 1.17 | 0.83 | |
| 01:04:29.41, +02:03:35.1 | 2016-12-21 / 01:16 | 1385 | 1.21 | 0.79 | |
| 01:04:29.39, +02:03:35.9 | 2016-12-21 / 01:41 | 1385 | 1.28 | 0.64 | |
| 01:04:29.43, +02:03:35.7 | 2016-12-22 / 01:05 | 1385 | 1.20 | 0.92 | |
| 01:04:29.44, +02:03:36.0 | 2016-12-22 / 01:30 | 1385 | 1.26 | 0.92 |